OpenStudy (anonymous):

Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. 5x+35y=-10 15x+5y=170

OpenStudy (anonymous):

5x+35y=-10 x=-7y-2 Substitute that into the second equation. 15(-7y-2)+5y=170 -105y-30+5y=170 -100y=200 y=-2 Then plug that back in. 5x+35(-2)=-10 5x-70=-10 5x=60 x=12

Parth (parthkohli):

Multiply first equation by 3 -> 15x + 105y = -30 15x = -30 - 105y Now, substitute 15x in the first equation with (-30 - 105y)

OpenStudy (anonymous):

ok